jabez177 (jabez177):

Ellie’s bird feeder is shaped like a cone with a height of 24 in. and a radius of 3 in. Packages of bird seed are cylindrical and come in different sizes. What is the height of the package that has a radius of 6 in. and contains exactly enough bird seed to fill the feeder? Use 3.14 to approximate pi. in.

jabez177 (jabez177):

@imqwerty

OpenStudy (haleyelizabeth2017):

I think you need to find the volume of the bird feeder first....

OpenStudy (haleyelizabeth2017):

\[V=\pi r^2 h\]

OpenStudy (haleyelizabeth2017):

So what is \[V=(3.14)(3)^2(24)?\]

jabez177 (jabez177):

678.24

OpenStudy (haleyelizabeth2017):

Right. Now: \[678.24=(3.14)(6)^2(x)\]

OpenStudy (haleyelizabeth2017):

Solve for x

jabez177 (jabez177):

You multiply those and then divide, right? :/

jabez177 (jabez177):

I got 6 for the height

OpenStudy (haleyelizabeth2017):

3.14*36 and yeah divide what you got for the volume for the bird feeder by what you get for 3.14*36

OpenStudy (haleyelizabeth2017):

That should be right, that is what I got

jabez177 (jabez177):

So it's 6? :/

OpenStudy (haleyelizabeth2017):

Yes, but I would have someone double check, just in case. Haven't done this for a while

jabez177 (jabez177):

@FortyTheRapper, can you check?

OpenStudy (fortytherapper):

I got 2 When you found the volume of a bird feeder, you used the same formula you did for the cylindrical figure You forgot to do the 1/3 part. The cone equals 226.08 after that, so \[226.08=(3.14)(6^2)(H)\]

jabez177 (jabez177):

Ohhh yeahhhh!!! The 1/3!!!!

OpenStudy (haleyelizabeth2017):

Oh...sorry D:

OpenStudy (fortytherapper):

Yeah, cone volume is the weirdest imo lol

jabez177 (jabez177):

yeah 2.000530 and on and on

jabez177 (jabez177):

so 2 is the answer?

OpenStudy (fortytherapper):

Yeah, after calculating I got \[226.08 = 113.04(h)\]Since pi = 3.14 was to be used

jabez177 (jabez177):

Yeah. Then I divided 226.08 by 113.04 And that got me 2

OpenStudy (fortytherapper):

Yep, sounds good

jabez177 (jabez177):

I have like 3 more.

OpenStudy (fortytherapper):

Aight, let's see em